# for Improving Reasoning

### v1

@@ 4.2 Setup for Evaluating Rationalyst
Evaluation tasks and metrics.
A summary of reasoning tasks we evaluate is provided in Table 2. We assess our method on the following datasets: GSM8K (Cobbe et al., 2021b) and MATH (Hendrycks et al., 2021) for mathematical reasoning, ECQA (Aggarwal et al., 2021) and HellaSwag (Zellers et al., 2019) for commonsense reasoning, ProofWriter (Tafjord et al., 2021) for logical reasoning, ARC (Clark et al., 2018) for scientific reasoning, and the recently proposed multi-task reasoning dataset MMLU-Pro (Wang et al., 2024) for holistic reasoning across multiple tasks. All tasks are evaluated using exact match as the metric. We apply the postprocessing setups from lm-evaluation-harness [^2] https://github.com/EleutherAI/lm-evaluation-harness before exact match calculation where applicable.
Task #Eval #Shots Reasoning Type
GSM8K 1319 8 Math
Math 5000 5 Math
ECQA 17944 6 CommonSense
HellaSwag 10000 4 CommonSense
ProofWriter 600 2 Logical
ARC 1172 4 Scientific
MMLU-Pro 12000 5 Mixed
Table 2: Configuration of the reasoning tasks tested. “#Eval” shows the number of instances used for evaluation. “#Shots” denotes the number of few-shot demonstrations provided for evaluation. For Math dataset, we report the results on all of the subsets. To evaluate ECQA, we use the validation split because the evaluation set requires running the evaluation on the original author’s server. For ProofWriter, we include only proofs with a depth of 5 or more. For ARC, we test on the ARC-Challenge subset.
Inference setting.
The model $M_{\text{Agent}}$ used for our baseline inference is also LLaMa-3-8B-Instruct. As mentioned earlier, to incorporate Rationalyst, we instruct $M_{\text{Agent}}$ to reason in a chain-of-thought manner. For procedural reasoning tasks like GSM8K, Math and MMLU_Pro, we provide in-context learning examples that break down the reasoning into individual steps leading to the final answer. For multiple-choice reasoning tasks like ECQA and ARC, we include examples that analyze and compare each answer choice. The content and number of in-context demonstrations align with lm-evaluation-harness or the original paper if available; otherwise, they are adjusted to fit the context window of LLaMa-3-8B-Instruct. Detailed prompts and in-context learning demonstrations are provided in Appendix B.
For all experiments, we employ a temperature of 0.7 during inference to facilitate sampling. This approach diverges from the conventional use of temperature 0, yielding improved performance on certain datasets (e.g., ProofWriter) while marginally reducing effectiveness on others (e.g., GSM8K). We set the sampling parameter $top\_k$ to 3, allowing $M_{\text{Agent}}$ to sample three reasoning steps simultaneously at each inference stage. For the baseline without Rationalyst, the next reasoning step is chosen randomly from these 3 samples. When using Rationalyst, the selection of the next step is guided by the rationales generated, as described in §3.3.
Other verifiers.
To evaluate the effectiveness of Rationalyst’s process supervision, we compare it with other approaches. For process supervision with other models, we include LLaMa-3-8B-Instruct and GPT-4 in our comparison. These models are prompted to rerank partial reasoning trajectories as reasoning steps are generated. The prompts and in-context learning demonstrations used for these models on representative datasets are provided in Appendix D. For outcome supervision, we also compare with outcome-based verifiers derived from LLaMa-3-8B-Instruct. These verifiers are fine-tuned on the training sets of each reasoning dataset. Following the approach outlined by Cobbe et al. (2021b), they assess the correctness of the final prediction by directly evaluating the question and final solution. This comparison allows us to assess the performance of Rationalyst against both process-based and outcome-based supervision methods.
@@ 5.1 Main result: Rationalyst Improves Performance on Various Tasks
Table 3: Accuracy and absolute improvement over baseline using Rationalyst. Rationalyst generalizes across different reasoning tasks, showing improved performance with unlabelled web-scale data.
@@ 5.2 Ablation: Web-scale Rationales Enhance Performance Across Tasks
Table 4: An ablation study on the benefit of rationales extracted from pre-training data (The Pile). We compare with Rationalyst against the “w/o Pile” model that is trained solely with rationales extracted from GSM8K and ECQA. The consistent accuracy drop shows that, utilizing web-scale rationales improves performance on various reasoning datasets.
@@ 5.4 Rationalyst Outperforms Other Verifiers
We observe consistent superior performance of Rationalyst compared to GPT-4’s process supervision. We hypothesize that this advantage arises from Rationalyst’s specialized design for providing supervision, in contrast to GPT-4’s general-purpose training.
@@
Table 6: Comparison of different supervision methods. Process supervision uses LLaMa-3 and GPT-4 to directly rerank each reasoning step. Outcome-based supervision fine-tunes LLaMa-3 on GSM8K and ECQA training sets to evaluate final answers. Rationalyst outperforms both strong verifiers like GPT-4 and similarly-sized models fine-tuned on matching training data.
@@ 5.5 Rationalyst Generates Accurate and Easy-to-understand Rationals
We annotate some samples from the test set of Math (Hendrycks et al., 2021) at inference time, which was not part of the rationale sampling datasets. Through manual observation, we find that our model can generate useful rationales that is helpful for understanding LLM’s reasoning process on Math (an example is provided in Appendix C). Comparing the rationales generated by Rationalyst with those generated by Quiet-Star (Zelikman et al., 2024) on the same problems, we find that our method produces more human-understandable rationales. We believe this happens because Quiet-Star optimizes rationales during training using the accuracy of the final prediction as a reward. This approach, while effective for improving task performance, does not explicitly prioritize human interpretability. In addition, this appraoch might inadvertently develop shortcuts or non-intuitive patterns that optimize for accuracy but not necessarily for clarity or human understanding.
@@ 6 Discussion
Scaling Rationalyst with stronger models and increased computational resources is a logical next step. Utilizing stronger models, such as LLaMa-3-70B or GPT-4, would enhance the quality of extracted rationales, improve filtration accuracy, and ultimately strengthen Rationalyst. However, due to computational constraints, we have not pursued this, which remains a limitation of this paper. Additionally, using larger unlabelled datasets with more extensive reasoning content, such as OpenWebMath (Paster et al., 2023), is currently infeasible due to the significant computational and time requirements for pre-filtering and training. These enhancements are planned for future work.
@@
Recent research has focused on extending computational resources at test-time (Snell et al., 2024; Wu et al., 2024), particularly for complex reasoning tasks. In our experiments, we focus on developing heuristics and employ a straightforward approach of sampling multiple candidates and reranking them based on Rationalyst’s guidance. However, Rationalyst’s framework is compatible with more sophisticated test-time compute techniques. Its heuristics can be integrated into existing algorithms like beam-search or look-ahead search, potentially enhancing their performance without significantly increasing computational cost.
@@
In our approach, we first select a subset of unlabelled data that contains strong reasoning signals, then extract implicit rationales from this data for model fine-tuning. While it has been demonstrated that training on data with robust reasoning signals can enhance reasoning capabilities on its own (Gunasekar et al., 2023; Jiang et al., 2024a), we believe our method offers additional performance benefits for two reasons. First, many language models have already been trained on datasets like The Pile. The value of fine-tuning on previously encountered text is likely lower than the value of fine-tuning on newly incorporated rationales. Second, implicit rationales encapsulate the reasoning process. Pre-training on these rationales enhances reasoning more effectively than focusing on the whole document.
@@ 7 Limitations
One limitation of this work is the comprehensiveness of our experiments. In future research, we plan to extend our experiments to a broader range of reasoning tasks and compare Rationalyst with other outcome-based and process-based verifiers. We also plan to adjust the combination of rationales used to train Rationalyst by (1) sampling from different reasoning tasks and (2) altering the mix of rationales in unlabelled web-scale pre-training data to better understand its generalizability.
8 Conclusion
In this paper, we introduced Rationalyst, a novel self-supervised model designed to enhance the reasoning capabilities of LLMs by leveraging hidden rationales extracted from unlabeled text. Our approach centers on the effective extraction and utilization of implicit rationales–those underlying thought processes that are not explicitly stated in the text but can be inferred. By capturing these rationales, Rationalyst provides a mechanism for process supervision during reasoning, enabling LLMs to reason better.
Acknowledgements.
We sincerely thank Eric Zelikman, Tianmin Shu, and the broader JHU CLSP community for discussions and inspiration.
Appendix A Prompts used for rationale sampling
@@
Appendix B Prompts used during inference
@@
Appendix C Examples of rationales generated at inference time
@@
Appendix D Prompts used for LLaMa-3 reranking

### v2

@@ 4.2 Setup for Evaluating Rationalyst
We evaluate Rationalyst across diverse reasoning tasks including mathematical (GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), commonsense (ECQA(Aggarwal et al., 2021) and HellaSwag (Zellers et al., 2019)), logical (ProofWriter (Tafjord et al., 2021)), scientific (ARC (Clark et al., 2018)), and multi-task reasoning (MMLU-Pro (Wang et al., 2024b)). For inference, we use LLaMa-3-8B-Instruct as our base agent model with temperature 0.7 and top-k sampling of 3. We compare Rationalyst against both process supervision approaches (using LLaMa-3-8B-Instruct and GPT-4) and outcome-based verifiers. Complete experimental details are provided in Appendix A.
4.3 Setup for other verifiers.
To evaluate the effectiveness of Rationalyst’s process supervision, we compare it with other approaches. For process supervision with other models, we include LLaMa-3-8B-Instruct and GPT-4 in our comparison. These models are prompted to rerank partial reasoning trajectories as reasoning steps are generated. The prompts and in-context learning demonstrations used for these models on representative datasets are provided in Appendix E. For outcome supervision, we also compare with outcome-based verifiers derived from LLaMa-3-8B-Instruct. These verifiers are fine-tuned on the training sets of each reasoning dataset. Following the approach outlined by Cobbe et al. (2021), they assess the correctness of the final prediction by directly evaluating the question and final solution. This comparison allows us to assess the performance of Rationalyst against both process-based and outcome-based supervision methods.
@@ 5.1 Main result: Rationalyst Improves Performance on Various Tasks
Table 2: Accuracy and absolute improvement over baseline using Rationalyst. Rationalyst generalizes improved performance across different reasoning tasks.
@@ 5.2 Ablation: Web-scale Rationales Enhance Performance Across Tasks
Table 3: An ablation study on the benefit of rationales extracted from pre-training data (The Pile). The consistent accuracy drop shows that, utilizing web-scale rationales improves performance on various reasoning datasets.
@@ 5.4 Rationalyst Outperforms Other Verifiers
Table 5: Comparison of different supervision methods. Rationalyst outperforms both strong verifiers like GPT-4 and similarly-sized models fine-tuned on matching training data.
From Table 5 we observe consistent superior performance of Rationalyst compared to GPT-4’s process supervision. We hypothesize that this advantage arises from Rationalyst’s specialized design for providing supervision, in contrast to GPT-4’s general-purpose training.
We would also like to mention that methods like self-consistency and other prompting techniques (e.g. STEP-BACK Prompting (Zheng et al., 2024)) are complementary to our approach. Rationalyst can be integrated with these methods to potentially achieve even better results. For more detail, please refer to Appendix H.
@@ 5.5 Rationalyst Generates Accurate and Easy-to-understand Rationals
We annotate some samples from the test set of Math (Hendrycks et al., 2021) at inference time, which was not part of the rationale sampling datasets. Through manual observation, we find that our model can generate useful rationales that is helpful for understanding LLM’s reasoning process on Math (an example is provided in Appendix D). Comparing the rationales generated by Rationalyst with those generated by Quiet-Star (Zelikman et al., 2024) on the same problems, we find that our method produces more human-understandable rationales. We believe this happens because Quiet-Star optimizes rationales during training using the accuracy of the final prediction as a reward. This approach, while effective for improving task performance, does not explicitly prioritize human interpretability. In addition, this appraoch might inadvertently develop shortcuts or non-intuitive patterns that optimize for accuracy but not necessarily for clarity or human understanding.
5.6 Rationalyst Introduces Minimal Inference Cost Overhead
We analyzed the inference costs of Rationalyst compared to baseline approaches. Using vLLM for efficient deployment with LLaMA-3-8B-Instruct as both the agent model and Rationalyst, we find minimal time overhead (within 5%) despite requiring two model calls per reasoning step, achieved through async batched execution and short rationale generation. Memory requirements approximately double with LLaMA-3-8B-Instruct, but this overhead becomes proportionally smaller with larger agent models (e.g., when using LLaMA-3-70B, Rationalyst’s additional 8B parameters represent a much smaller relative increase). These computational costs are minimal compared to Rationalyst’s demonstrated accuracy improvements, making the method efficient for applications prioritizing reasoning quality.
@@ 6 Discussion
Scaling Rationalyst with stronger models and increased computational resources is a logical next step. Utilizing stronger models, such as LLaMa-3-70B or GPT-4, would enhance the quality of extracted rationales, improve filtration accuracy, and ultimately strengthen Rationalyst. However, due to computational constraints, we have not pursued this. Additionally, using larger unlabelled datasets with more extensive reasoning content, such as OpenWebMath (Paster et al., 2023), is currently infeasible due to the significant computational and time requirements for pre-filtering and training.
@@
Recent research has focused on extending computational resources at test-time (Snell et al., 2024; Wu et al., 2024), particularly for complex reasoning tasks. In our experiments, we focus on developing heuristics and employ a straightforward approach of sampling multiple candidates and reranking them based on Rationalyst’s guidance. However, Rationalyst’s framework is compatible with more sophisticated test-time compute techniques. Its heuristics can be integrated into existing algorithms like beam-search or look-ahead search.
@@
While it has been demonstrated that training on data with robust reasoning signals can enhance reasoning capabilities on its own (Gunasekar et al., 2023; Jiang et al., 2024a), we believe our method offers additional performance benefits for two reasons. First, many language models have already been trained on datasets like The Pile. The value of fine-tuning on previously encountered text is likely lower than the value of fine-tuning on newly incorporated rationales. Second, implicit rationales encapsulate the reasoning process. Pre-training on these rationales enhances reasoning more effectively than focusing on the whole document.
7 Conclusion and Future Work
In this paper, we introduced Rationalyst, a novel self-supervised model designed to enhance the reasoning capabilities of LLMs by leveraging hidden rationales extracted from unlabeled text. While our results demonstrate the effectiveness of this approach, there are several promising directions for future work, including scaling to larger models and datasets, integrating with test-time compute optimization techniques, and exploring the role of implicit rationales in model training. We hope these insights will inspire further research into leveraging latent knowledge in pre-training data to enhance model capabilities.
Acknowledgments
This work is supported by ONR grant (N00014-241-2089). The GPUs were provided by the DSAI cluster. We sincerely thank Eric Zelikman, Tianmin Shu, and the broader JHU CLSP community for discussions and inspiration.
@@ 8 Limitations
One limitation of this work is the comprehensiveness of our experiments. In future research, we plan to extend our experiments to a broader range of reasoning tasks and compare Rationalyst with other outcome-based and process-based verifiers. We also plan to adjust the combination of rationales used to train Rationalyst by (1) sampling from different reasoning tasks and (2) altering the mix of rationales in unlabelled web-scale pre-training data to better understand its generalizability. An additional limitation is that we could potentially improve performance through preference tuning (e.g., DPO) where the model learns to distinguish between valid and invalid rationales, but we have not explored this direction.
Appendix A Evaluation Setup Details
A.1 Task Configuration and Metrics
We evaluate Rationalyst on a diverse set of reasoning tasks, carefully selected to assess performance across different reasoning domains. Table 6 provides a comprehensive overview of our evaluation configuration.
Task #Eval #Shots Reasoning Type
GSM8K 1319 8 Math
Math 5000 5 Math
ECQA 17944 6 CommonSense
HellaSwag 10000 4 CommonSense
ProofWriter 600 2 Logical
ARC 1172 4 Scientific
MMLU-Pro 12000 5 Mixed
Table 6: Detailed configuration of evaluation tasks. For each dataset, we report the training set size (#Train), evaluation set size (#Eval), and number of few-shot demonstrations (#Shots) used during evaluation. We use exact match accuracy across all tasks, following standard evaluation protocols.
For each dataset, we implement specific evaluation protocols:
  • GSM8K: We evaluate on the standard test split, using 8-shot demonstrations that showcase step-by-step mathematical reasoning.
  • MATH: Our evaluation encompasses all difficulty levels and mathematical topics in the dataset, utilizing 5-shot demonstrations calibrated for complexity.
  • ECQA: We use the validation split due to evaluation server constraints for the test set, maintaining consistency with prior work.
  • HellaSwag: Evaluation follows the original setup with 4-shot demonstrations focusing on commonsense inference.
  • ProofWriter: We specifically evaluate on proofs with depth more than 5 to assess complex logical reasoning capabilities.
  • ARC: Our evaluation focuses on the more challenging ARC-Challenge subset, employing 4-shot demonstrations.
  • MMLU-Pro: Following the original paper’s protocol, we evaluate across all reasoning categories using 5-shot demonstrations.
A.2 Inference Configuration
Our inference setup utilizes LLaMa-3-8B-Instruct as the base agent model ($M_{\text{Agent}}$). Key configuration parameters include:
Sampling Parameters.
We employ temperature = 0.7 and top-k = 3 during inference. This configuration allows for diverse reasoning path exploration while maintaining coherence. The temperature setting represents a calculated trade-off between deterministic output (temperature = 0) and exploration, showing particular effectiveness on logical reasoning tasks like ProofWriter.
Chain-of-Thought Implementation.
For each task category, we implement specialized chain-of-thought prompting:
  • Mathematical and Procedural Tasks: Demonstrations break down solutions into atomic reasoning steps, with explicit intermediate calculations.
  • Multiple-Choice Tasks: Prompts guide systematic analysis of each answer choice, encouraging comparative reasoning.
  • Logical Reasoning: Step-by-step deductive reasoning chains are demonstrated, emphasizing premise utilization.
Appendix B Prompts used for rationale sampling
@@
Appendix C Prompts used during inference
@@
Appendix D Examples of rationales generated at inference time
@@
Appendix E Prompts used for LLaMa-3 reranking
@@
Appendix F Extended Related Work
Supervising reasoning.
Supervision-based approaches have been shown to enhance the reasoning abilities of LLMs. Cobbe et al. (2021) and Snell et al. (2024) demonstrate that training a “verifier" to supervise reasoning can be more parameter-efficient than simply expanding the parameters of the “reasoner" responsible for solving the reasoning task. Ground-truth feedback from interaction with the environment is an effective form of supervision (Wang et al., 2023), but it works only in controlled environments like simulated world. Some verifiers employ principles like compositional reasoning to validate the reasoning process (Dhuliawala et al., 2023; Weir et al., 2024, 2023; Vacareanu et al., 2024), these general-purpose approaches do not fully leverage the vast amount of available unlabeled data that a data-driven approach could utilize. Process-based supervision (Lightman et al., 2023; Luo et al., 2024; Wang et al., 2024a) offers supervision at each reasoning step rather than just at the final result. While promising, it requires substantial human annotation for the correctness of intermediate steps or the existence of ground-truth answers for data collection. Our work aims to address these challenges by proposing a data-centric process-supervision method without the need for human annotation.
Knowledge extraction from unlabelled data.
Rationales as the basis for reasoning.
Appendix G The Resulting Data from Extraction/Filteration.
On GSM8K, our method generates an average of 2.34 rationales per document, while on ECQA, it generates 2.58 rationales per document. The filtration process removes 80.5% of the generated rationales on GSM8K and 42.4% on ECQA.
Appendix H Rationalyst in Combination with General Methods for Reasoning Enhancement
We have conducted additional experiments focusing on GSM8K and Math datasets to verify our claim. We specifically chose these arithmetic reasoning benchmarks over multiple-choice questions, as using self-consistency on multiple-choice tasks could artificially inflate performance. Using 64 generations for self-consistency, we observed improved accuracy on both benchmarks with the help of Rationalyst.